OpenStudy (anonymous):

Events E and F in a sample space S. Assume that Pr[E]=17/30, Pr[F|E]=12/17, and Pr[E′∩F′]=1/5. (1) Pr[F]= (2) Pr[E∩F]=

OpenStudy (anonymous):

#2 is 102/255

OpenStudy (sleepyjess):

Welcome to OpenStudy! I'm not sure on this, but maybe @freckles or @hero can help :)

OpenStudy (freckles):

what!?! what is PR mean? Please don't say probability.

OpenStudy (anonymous):

question makes no sense to me, a probability is never larger than 1

OpenStudy (anonymous):

sorry i fixed it @satellite73

OpenStudy (anonymous):

that's better

OpenStudy (anonymous):

\[P(F|E)=\frac{P(F\cap E)}{P(E)}\]

OpenStudy (anonymous):

you can use that to find \[P(F\cap E)\]

OpenStudy (anonymous):

i already found F∩E which is 102/255

OpenStudy (anonymous):

oh no !

OpenStudy (anonymous):

\[P(F\cap E)=\frac{17}{30}\times \frac{12}{17}=\frac{12}{30}=\frac{2}{5}\]

OpenStudy (pulsified333):

@satellite73 isn't P(F∩E) = Pr(F|E) * Pr(E)?